OpenStudy (anonymous):

Solve for x. Rs/x =1/4 A. X=rs/4 B. X=4/rs C. X=1/4rs D. X=4rs

geerky42 (geerky42):

\[\dfrac{rs}{x}=\dfrac{1}{4}~~~\Longrightarrow~~~\dfrac{x}{rs}=4\]Right? Now multiply both sides by \(rs\).

OpenStudy (anonymous):

I got its D?